OpenStudy (anonymous):

RUSH! The sum of finite geometric series 1+5+25+125+...+5^9 answer please???

OpenStudy (anonymous):

a1=1,r=5 sum = a1(1-r^n)/1-r

OpenStudy (anonymous):

here a1=1 and r=5 and n=10 use the formula for summation of GP with r>1 Sn=1(5^10 -1) / (5-1)=1/4 * (5^10 -1)

OpenStudy (anonymous):

I TRy but the answer i got is 488281. something : l

OpenStudy (anonymous):

S = a((1-r^n) / (1-r)) = 1(1-5^10) / (1-5) = 2,441,406

OpenStudy (anonymous):

thank u ;)